A chemist has 0.2 liters of an acid solution whose acid concentration is 65%. How many liters of pure water should be added to lower the concentration to 50%?

Answer by robertb(5830)

Let w = liters of water to add.
==>0.65*0.2 + 0.00*w = 0.50*(0.2+w)
==> 0.13 = 0.1 + 0.5w ==> 0.03 = 0.5w ==> w = 0.06 Liters of water should be added.